### v1

@@ Abstract
Spectrally truncated three-dimensional (3D) Euler flows reach an absolute equilibrium. Yet they show non-trivial temporal correlations, similar to the ones of turbulent flows. In this work, we investigate Navier-Stokes equation in equilibrium conditions, both at finite viscosity and in the inviscid limit. We determine the space-time velocity correlations, using functional renormalisation group and direct numerical simulations. We characterise the crossover in the dynamical scaling from $\tau\sim 1/(\nu k^{2})$ at finite viscosity to $\tau\sim 1/(u_{\rm rms}k)$ in the inviscid limit, where $\tau$ is the decorrelation time. We show that, although the static properties of spectrally truncated Euler flows radically differ from their hydrodynamical limit, they share a similar emergent $k^{-1}$ dynamical scaling.
@@ 1 Introduction
Since the pioneering works of Kolmogorov in 1941, deriving the statistical theory of turbulence from the fundamental hydrodynamical equations has remained an unsolved challenge. To gain some understanding, simple models, such as the three-dimensional (3D) Euler equation or the one-dimensional (1D) Burgers equation, have played a pivotal role. A noticeable feature of these models is that, while the continuous equations may generate singularities in finite time [1, 2, 3], their spectrally (Galerkin)-truncated versions lead to thermalisation to an absolute equilibrium state [4, 5, 6]. The static properties of the continuous and spectrally-truncated models thus radically differ. For the 3D Euler flow, the continuous equation is expected to yield a turbulent non-equilibrium state with Kolmogorov kinetic energy spectrum $E(k)\sim k^{-5/3}$, whereas the spectrally-truncated equation leads to the equilibrium state characterised by an equipartition energy spectrum $E(k)\sim k^{d-1}$. Note that the equipartition spectrum can also be observed in turbulent Navier-Stokes flows in the presence of thermal noise at sufficiently small scales in the dissipative range [7].
@@
The thermalisation process to absolute equilibrium has been thoroughly characterised, and was shown to involve localised oscillatory structures called tygers, both for the 1D Burgers and for the 3D Euler spectrally-truncated equations [8, 9, 10, 11]. The emergence of these tygers deeply imprints the dynamics, yielding a relaxation time $\tau(k)\sim k^{-z}$ with $z=1$, both in the 1D Burgers [8] and in the 3D Euler [12, 13] equations. Interestingly, when a thermal noise is added to the deterministic 1D Burgers equation, the latter maps to the Kardar-Parisi-Zhang (KPZ) equation [14], a celebrated Langevin-type model in statistical physics which features non-equilibrium scaling phenomena [15, 16]. It was shown in direct numerical simulations that the stochastic 1D Burgers–KPZ equation also exhibits a decorrelation time scaling $\tau(k)\sim k^{-1}$ in the inviscid limit, as its deterministic counterpart [17, 18].
@@
This finding weaves a deep connection between statistical physics and deterministic hydrodynamics. It turns out that this link has not been explored yet for the spectrally-truncated 3D Euler equation. The stochastic version of the 3D Navier-Stokes equation in the presence of thermal noise was introduced in the seminal paper by Forster, Nelson, Stephen (FNS) [19, 20], and named model A. The authors showed that this equation also leads to relaxation to an absolute equilibrium state, and thus model A corresponds to a Langevin model for a fluid near thermal equilibrium. Let us notice that such studies of near-equilibrium (compressible) fluids date back to the early works by Landau and Placzek [21, 22] and have important applications. However, the stochastic Euler dynamics was not discussed by FNS. In this paper, we provide the missing link, by investigating the inviscid limit of model A, using both numerical simulations and functional renormalisation group (FRG).
@@
For the numerical part, we perform direct numerical simulations of model A of FNS, ie we integrate the spectrally-truncated stochastic 3D Navier-Stokes equation, in a Taylor-Green configuration, which allows us to access large system sizes. We compute the spatio-temporal two-point correlation function for different microscopic viscosities to determine the behaviour of the half decay time $\tau_{1/2}(k)$. We find that it also displays the $z=1$ scaling in the inviscid limit, as for the deterministic 3D Euler dynamics.
@@
To present the FRG part, let us first summarise previous achievements on which this work builds. For the stochastic 1D Burgers–KPZ equation, the origin of the $z=1$ scaling has been recently elucidated using FRG [23]. Indeed, it was shown to be controlled by a new fixed point of the Burgers–KPZ equation termed inviscid Burgers (IB) fixed point, and characterised by a critical dynamical exponent $z=1$, differing from the well-known KPZ value $z=3/2$. The FRG is a field-theoretical method based on the continuous equations. This salient scaling of the decorrelation time is not affected by the spectral truncation, although the static properties are drastically altered. Let us notice that for the 3D Navier-Stokes equation in the turbulent state, the emergence of a $z=1$ scaling is well-known phenomenologically, and interpreted as a consequence of random sweeping effect [24, 25, 26, 27, 28, 29, 30, 31]. For the forced Navier-Stokes, a rigorous derivation of this $z=1$ was provided using FRG [33, 34, 35].
@@
In this work, we use FRG to revisit model A of FNS, focusing in particular on its inviscid limit. We show that, beside the attractive (IR) fixed point identified by FNS, there exists another repulsive (UV) fixed point, associated with zero viscosity, which we call inviscid FNS (IFNS), in analogy with the IB fixed point for the stochastic Burgers–KPZ equation, and which is also characterised by a $z=1$ critical dynamical exponent. Thus, for both the stochastic 1D Burgers and 3D Navier-Stokes equations in equilibrium conditions, a non-trivial behaviour emerges in the inviscid limit, which is distinct from the KPZ or FNS one. Moreover, we show that this behaviour is also relevant in the viscous case, where the corresponding scaling is observable at large momentum, over a range that extends when the viscosity is further reduced.
@@
The remainder of the paper is organised as follows. In Sec. 2, we introduce model A of FNS, and review the main results obtained by FNS relevant for this work. In Sec. 3, we briefly present the FRG formalism, and derive the flow equations for model A. We establish the fixed point structure of model A, and show that the dynamical exponent at the IFNS fixed point is $z=1$, both through a numerical integration of the complete flow equations within some approximation, and from an analytic solution of exact asymptotic flow equations in the large momentum limit. In Sec. 4, we present the results of the direct numerical simulations of model A, and compare with the FRG results. Finally, our conclusion is given in Sec. 5.
@@ 2.1 Model A of FNS
where $\mathbf{v}(t,\mathbf{x})$ is the fluid velocity field, $\pi(t,\mathbf{x})$ the pressure field, $\nu$ the (microscopic) kinematic viscosity of the fluid, $\rho$ its density, and $\lambda$ a real parameter introduced for convenience and eventually set to unity. The field $\mathbf{f}$ is a stochastic forcing with Gaussian distribution of zero mean and of two-point correlations given in Fourier space by
@@
Because of translational invariance in space and time, the total frequency and momentum are conserved. Thus, the Fourier transform of a generic $n$-point function depends on only $n-1$ frequencies and momenta since the last ones are fixed by conservation. We denote with an overbar the Fourier transform without the factor $(2\pi)^{d+1}\delta\left(\sum_{k=1}^{n}\omega_{k}\right)\delta^{d}\left(\sum_{k=1}^{n}\mathbf{q}_{k}\right)$ stemming from the conservation laws.
The choice $\bar{D}(q)=Dq^{2}$ models a thermal noise, and the set of equations (1)–(3) identifies in this case with model A of FNS [19, 20]. It corresponds to a Langevin dynamics for a fluid near thermal equilibrium. In particular, it satisfies a fluctuation-dissipation theorem which fixes the amplitude of the noise to $D=\frac{\nu}{\rho}k_{B}T$ where $T$ is the temperature of the fluid and $k_{B}$ the Boltzmann constant. This system relaxes at long times to an absolute equilibrium state, with equipartition of energy, yielding a kinetic energy spectrum $\bar{E}(k)\sim k^{d-1}$. In this state, the velocity has a Gaussian distribution and its two-point correlation function takes the scaling form
@@
The constitutive equations of model A can be rescaled, such that there remains a single parameter $\bar{\lambda}=\lambda D^{1/2}/{\nu}^{3/2}$. As emphasised by FNS [20], the absolute equilibrium state corresponds to a vanishing renormalised $\bar{\lambda}$. This is why the correlations (6)–(7) are similar to results from linearised hydrodynamics. Moreover, the near equilibrium dynamics can be captured using perturbation theory, which is organised as a power series in $\bar{\lambda}$. However, this approach is not suitable to study fully developed turbulence, which corresponds to a non-trivial steady-state with a finite $\bar{\lambda}$ and non-Gaussian fluctuations. The perturbation theory in $\bar{\lambda}$ obviously also fails to study the inviscid limit $\nu\to 0$, even in the equilibrium state. For this, one has to resort to non-pertubative methods, such as the FRG which we explain in Sec. 3.
@@ 2.2 Action and symmetries for model A
In order to apply renormalisation group techniques, one can cast the stochastic equation into a path integral formulation, following the Martin-Siggia-Rose-Janssen-de Dominicis formalism [37, 38, 39]. One obtains from the stochastic Navier-Stokes equation with thermal noise (1)–(3) the following generating functional and action [35]
@@
where $\bar{\mathbf{v}}$ and $\bar{\pi}$ are the response velocity and response pressure fields, and $\mathbf{j},\bar{\mathbf{j}},K,\bar{K}$ are the sources linearly coupled to the fields. The functional ${\cal W}=\ln{\cal Z}$ is the generating functional of connected correlation functions – which are the equivalent, for fluctuating fields instead of random variables, to cumulants. Indeed, the connected correlation and response functions are obtained as
@@
$\Gamma[\Psi]=\underset{\mathcal{J}}{\sup}\left[-\mathcal{W}[\mathcal{J}]+\int_{t,\mathbf{x}}\mathcal{J}\Psi\right]\,,\qquad\hbox{where}\quad\Psi_{i}\equiv\langle\Phi_{i}\rangle_{\cal J}=\dfrac{\delta{\cal W}}{\delta{\cal J}_{i}}\,,$ (12)
with $\Phi=(\mathbf{v},\bar{\mathbf{v}},\pi,\bar{\pi})$ denoting the multiplet of fields, $\mathcal{J}=(\mathbf{j},\bar{\mathbf{j}},K,\bar{K})$ the associated sources, and $\Psi=(\mathbf{u},\bar{\mathbf{u}},\Pi,\bar{\Pi})$ the expectation values of the fields in the presence of the sources ${\cal J}$. One can obtain from $\Gamma$ the vertex functions, defined as
$\Gamma^{(m,n)}_{\alpha_{1}...\alpha_{m+n}}(t_{1},\mathbf{x}_{1};\ldots;t_{m+n},\mathbf{x}_{m+n})\equiv\dfrac{\delta^{m+n}\Gamma}{\delta u_{\alpha_{1}}(t_{1},\mathbf{x}_{1})\ldots\delta\bar{u}_{\alpha_{m+n}}(t_{m+n},\mathbf{x}_{m+n})}\Big{|}_{\mathbf{u},\bar{\mathbf{u}}=0}\,.$ (13)
The knowledge of the set of $G^{(m,n)}$, or equivalently of the set of $\Gamma^{(m,n)}$, provides the statistical theory for the model under study.
Let us now briefly explain how the symmetries lead to exact relations between these functions, which are called Ward identities, and which can be very efficiently exploited within the FRG formalism. In fact, it turns out to be particularly useful to consider extended symmetries, which are continuous infinitesimal transformations which do not leave the action strictly invariant, but yields a variation at most linear in the fields. One can indeed derive from such extended symmetries exact Ward identities, which deeply constrain the set of $G^{(m,n)}$ (or equivalently $\Gamma^{(m,n)}$[40].
This procedure has been expounded in details, in eg. Refs. [34, 35] for the Navier-Stokes equation, or in eg. Refs. [41, 23] for the related Burgers–KPZ equation. Hence, we only provide here a summary, and refer the interested reader to these references for technical details. The first noticeable extended symmetries are shifts of the pressure fields, $\pi(t,\mathbf{x})\to\pi(t,\mathbf{x})+\epsilon(t,\mathbf{x})$, and $\bar{\pi}(t,\mathbf{x})\to\bar{\pi}(t,\mathbf{x})+\bar{\epsilon}(t,\mathbf{x})$, with $\epsilon(t,\mathbf{x})$ and $\bar{\epsilon}(t,\mathbf{x})$ infinitesimal scalar functions of time and coordinate. The corresponding Ward identities impose that the variation of the effective action $\Gamma$ under these transformations is equal to that of the action ${\cal S}$, which entails that the whole pressure sector is not renormalised. A second extended symmetry is the time-dependent Galilean symmetry, which corresponds to the transformation
@@
with $\epsilon(t)$ an infinitesimal vectorial function of time. The corresponding Ward identities exactly relate a vertex function $\Gamma^{(m+1,n)}$ with one vanishing velocity momentum to lower order ones $\Gamma^{(m,n)}$. We here simply give the first two identities
@@
with $\bar{\epsilon}(t)$ an infinitesimal vectorial function of time. The corresponding Ward identities exactly impose that a vertex function $\Gamma^{(m,n)}$ with one vanishing response velocity momentum vanish, except the two lowest order ones, which are fixed to
@@
This symmetry also yield constraints on the vertex functions, in particular

### v2

@@ Abstract
In this work, we investigate the Navier-Stokes equation in the presence of thermal noise, both at finite viscosity (revisiting the seminal work by Forster-Nelson-Stephen) and in the inviscid limit, which has not yet been explored. We determine the space-time velocity correlations in this dynamics, using functional renormalisation group and direct numerical simulations. While spectrally truncated three-dimensional Euler flows reach a stationary equilibrium state, they exhibit non-trivial temporal correlations. We show that these non-trivial correlations persist for small but finite viscosity, yielding an emergent $\tau\sim k^{-1}$ dynamical scaling, where $\tau$ is the decorrelation time. We characterise the crossover from the scaling $\tau\sim 1/(\nu k^{2})$, expected at large viscosity, to the scaling $\tau\sim 1/(u_{\rm rms}k)$ found in the inviscid limit.
keywords:
stochastic Navier-Stokes equation, renormalisation group, direct numerical simulations
{fmtext}
@@ 1 Introduction
Deriving the statistical theory of turbulence from the fundamental hydrodynamical equations has remained an unsolved challenge, despite intensive efforts since the pioneering works of Kolmogorov in 1941. To gain some understanding, models such as the three-dimensional (3D) Euler equation or the one-dimensional (1D) Burgers equation, have played a pivotal role. A noticeable feature of these models is that, while the continuous equations may generate singularities in finite time [1, 2, 3, 4], their
spectrally (Galerkin)-truncated versions lead to thermalisation to an equilibrium state [5, 6, 7, 8, 9, 10, 11, 12, 13]. The static properties of the continuous and spectrally-truncated models thus radically differ. For the 3D Euler flow, the continuous equation is expected to yield a turbulent non-equilibrium state with Kolmogorov kinetic energy spectrum $E(k)\sim k^{-5/3}$, whereas the spectrally-truncated equation leads to the equilibrium state characterised by an equipartition energy spectrum $E(k)\sim k^{2}$. Note that the equipartition spectrum can also be observed in turbulent Navier-Stokes flows in the presence of thermal noise at sufficiently small scales in the dissipative range [14, 15, 16].
@@
The thermalisation process to equilibrium has been thoroughly characterised, and was shown to involve localised oscillatory structures called tygers, both for the 1D Burgers and for the 3D Euler spectrally-truncated equations [7, 8, 9, 10]. The emergence of these tygers deeply imprints the dynamics, yielding a relaxation time $\tau(k)\sim k^{-z}$ with $z=1$, both in the 1D Burgers [7] and in the 3D Euler [11, 13] equations. Interestingly, when a thermal noise is added to the deterministic 1D Burgers equation, the latter maps to the Kardar-Parisi-Zhang (KPZ) equation [17], a celebrated Langevin-type model in statistical physics which features non-equilibrium scaling phenomena [18, 19]. It was shown in direct numerical simulations that the stochastic 1D Burgers–KPZ equation also exhibits a decorrelation time scaling $\tau(k)\sim k^{-1}$ in the inviscid limit, as its deterministic counterpart [20, 21].
@@
This finding weaves a deep connection between statistical physics and deterministic hydrodynamics. It turns out that this link has not been explored yet for the spectrally-truncated 3D Euler equation. The aim of this paper is to bridge this gap, by investigating the stochastic Navier-Stokes equation in the presence of a thermal noise, and uncovering what happens when taking the inviscid limit of this model. The stochastic version of the 3D Navier-Stokes equation, including thermal noise, was introduced in the seminal paper by Forster, Nelson, Stephen (FNS) [22, 23], and named model A. The authors showed that this equation also leads to relaxation to an equilibrium state, and thus model A corresponds to a Langevin model for a fluid near thermal equilibrium. Let us notice that such studies of near-equilibrium (compressible) fluids date back to the early works by Landau and Placzek [24, 25] and have important applications. However, the stochastic Euler dynamics was not discussed by FNS, as their approach relies on perturbative dynamical RG, which does not allow one to access the non-perturbative inviscid limit.
Beyond the theoretical interest, let us emphasise the physical motivation to revisit model A. Let us remind that the hydrodynamical description is an effective description valid down to some length scale, which we denote $\Lambda^{-1}$. Hence, one has to account for fluctuations coming from the molecular dynamics below this scale. Adding a thermal noise, thereby yielding the stochastic Navier-Stokes equation, is therefore physically relevant (see for example Refs. [14, 26] and references within). The work [14] highlights the appearance of the thermal noise as an observable effect, at lengthscales where the fluctuating hydrodynamics is still valid (see Eq. (2.14) in [14] and the discussion after). This was confirmed by numerical simulations [15]. In fact, the observation of an equipartition spectrum $E(k)\propto k^{2}$ at lengthscales comparable to the Kolmogorov scale was first reported in incompressible fluid turbulence experiments [16]. In this (dissipative) range, the turbulent energy spectrum decays exponentially [27], such that the thermal fluctuations dominate over the turbulent ones. This provides a further motivation to include the thermal noise in the Navier-Stokes equation.
@@
In this work, we thus focus on model A, and study its yet unexplored inviscid limit, using both numerical simulations and functional renormalisation group (FRG). For the numerical part, we perform direct numerical simulations of model A of FNS, ie we integrate the spectrally-truncated stochastic 3D Navier-Stokes equation, in a Taylor-Green configuration, which allows us to access large system sizes. We compute the spatio-temporal two-point correlation function for different microscopic viscosities to determine the behaviour of the half decay time $\tau_{1/2}(k)$. We find that it also displays the $z=1$ scaling in the inviscid limit, as for the deterministic 3D Euler dynamics.
@@
To present the FRG part, let us first summarise previous achievements on which this work builds. For the stochastic 1D Burgers–KPZ equation, the origin of the $z=1$ scaling has been recently elucidated using FRG [28]. Indeed, it was shown to be controlled by a new fixed point of the Burgers–KPZ equation termed inviscid Burgers (IB) fixed point, and characterised by a critical dynamical exponent $z=1$, differing from the well-known KPZ value $z=3/2$. The FRG is a field-theoretical method based on the continuous equations. This salient scaling of the decorrelation time is not affected by the spectral truncation, although the static properties are drastically altered. Let us notice that for the 3D Navier-Stokes equation in the turbulent state, the emergence of a $z=1$ scaling is well-known phenomenologically, and interpreted as a consequence of random sweeping effect [29, 30, 31, 32, 33, 34, 35, 36]. For the forced Navier-Stokes, a rigorous derivation of this $z=1$ was provided using FRG [37, 38, 39]. Let us notice that it has been proposed in [40], and used in several subsequent works, to move from the Eulerian to a quasi-Lagrangian frame in order to get rid of the sweeping effect, which hinders the perturbative approaches [30]. However, within the FRG formalism, which is non-perturbative, the sweeping effect is not problematic, and does not prevent from correctly describing the subleading straining turbulent dynamics and the Kolmogorov spectrum [39]. Thus, it is not necessary in this formalism to resort to a quasi-Lagrangian description, which is more complicated than the Eulerian one.
@@
Here, we use FRG to revisit model A of FNS, focusing in particular on its inviscid limit. We show that, beside the attractive (IR) fixed point identified by FNS, there exists another repulsive (UV) fixed point, associated with zero viscosity, which we call inviscid FNS (IFNS), in analogy with the IB fixed point for the stochastic Burgers–KPZ equation, and which is also characterised by a $z=1$ critical dynamical exponent. Thus, for both the stochastic 1D Burgers and 3D Navier-Stokes equations in equilibrium conditions, a non-trivial behaviour emerges in the inviscid limit, which is distinct from the KPZ or FNS one. Moreover, we show that this behaviour is also relevant in the viscous case, where the corresponding scaling is observable at large momentum, over a range that extends when the viscosity is further reduced.
@@
The remainder of the paper is organised as follows. In Sec. 2, we introduce model A of FNS, and review the main results obtained by FNS which are relevant for this work. In Sec. 3, we briefly present the FRG formalism, and derive the flow equations for model A. We establish the fixed point structure of model A, and show that the dynamical exponent at the IFNS fixed point is $z=1$, both through a numerical integration of the complete flow equations within some approximation, and from an analytic solution of exact asymptotic flow equations in the large momentum limit. In Sec. 4, we present the results of the direct numerical simulations of model A, and compare with the FRG results. Finally, we discuss the physical relevance of the IFNS scaling in Sec. 5 before giving our conclusions in Sec. 6.
@@ 2.1 Model A of FNS
where $\mathbf{v}(t,\mathbf{x})$ is the fluid velocity field, $\pi(t,\mathbf{x})$ the pressure field, $\nu$ the (microscopic) kinematic viscosity of the fluid, $\rho$ its density, and $\lambda$ a real parameter introduced for later purpose and eventually set to unity. The field $\mathbf{f}$ is a stochastic forcing with Gaussian distribution of zero mean and of two-point correlations given in Fourier space by
@@
We consider an infinite volume fluid, which we assume to be homogeneous and isotropic, and we focus on the stationary state. The system is thus translationally invariant in space and time, and as a consequence the total frequency and momentum are conserved. Thus, the Fourier transform of a generic $n$-point function depends on only n1n-1 frequencies and momenta since the last ones are fixed by conservation. We denote with an overbar the Fourier transform without the factor $(2\pi)^{d+1}\delta\left(\sum_{k=1}^{n}\omega_{k}\right)\delta^{d}\left(\sum_{k=1}^{n}\mathbf{q}_{k}\right)$ stemming from the conservation laws.
The choice $\bar{D}(q)=Dq^{2}$ models a thermal noise, and the set of equations (1)–(3) identifies in this case with model A of FNS [22, 23]. It corresponds to a Langevin dynamics for a fluid near thermal equilibrium. In particular, it satisfies a fluctuation-dissipation theorem which fixes the amplitude of the noise to $D=\frac{\nu}{\rho}k_{B}T$ where $T$ is the temperature of the fluid and $k_{B}$ the Boltzmann constant. This system relaxes at long time to a stationary equilibrium state, with equipartition of energy, yielding a kinetic energy spectrum $\bar{E}(k)\sim k^{d-1}$. This stochastic dynamics is invariant under the time-reversal symmetry defined by the tranformation (18) (see next section). For this reason, we refer to the stochastic Navier-Stokes with thermal noise (model A) as the ‘equilibrium’ dynamics, as it leads to an equilibrium stationary state (fluid at rest), while the Navier-Stokes equation with forcing at large scale is referred to as ‘non-equilibrium’ since it leads to a stationary turbulent state. The forcing at large scales indeed breaks the symmetry (18) and the fluctuation-dissipation relation no longer holds in the turbulent state  [^1] Note that in this work, equilibrium does not refer to the Euler dynamics and non-equilibrium to the Navier-Stokes dynamics, which is another commonly used definition..
In this state, the velocity has a Gaussian distribution and its two-point correlation function takes the scaling form
@@
The constitutive equations of model A can be rescaled, such that there remains a single parameter $\bar{\lambda}=\lambda D^{1/2}/{\nu}^{3/2}$. As emphasised by FNS [23], the equilibrium state corresponds to a vanishing renormalised $\bar{\lambda}$. This is why the correlations (6)–(7) are similar to results from linearised hydrodynamics. Moreover, the near equilibrium dynamics can be captured using perturbation theory, which is organised as a power series in $\bar{\lambda}$. However, this approach is not suitable to study fully developed turbulence, which corresponds to a non-trivial steady-state with a finite $\bar{\lambda}$ and non-Gaussian fluctuations. The perturbation theory in $\bar{\lambda}$ obviously also fails to study, even in the equilibrium state, the inviscid limit $\nu\to 0$, as it corresponds to $\bar{\lambda}\to\infty$. For this, one has to resort to non-pertubative methods, such as the FRG which we explain in Sec. 3.
@@ 2.2 Action and symmetries for model A
In order to apply renormalisation group techniques, one can cast the stochastic dynamical equation into a path integral formulation, following the Martin-Siggia-Rose-Janssen-de Dominicis formalism [42, 43, 44]. One obtains from the stochastic Navier-Stokes equation with thermal noise (1)–(3) the following generating functional and action [39]
@@
where $\bar{\mathbf{v}}$ and $\bar{\pi}$ are the response velocity and response pressure fields, and $\mathbf{j},\bar{\mathbf{j}},K,\bar{K}$ are the sources linearly coupled to the fields. The functional ${\cal W}=\ln{\cal Z}$ is the generating functional of connected correlation functions – which are the equivalent, for fluctuating fields instead of random variables, to cumulants. Indeed, the connected velocity correlation and response functions are obtained as
@@
$\Gamma[\Psi]=\underset{\mathcal{J}}{\sup}\left[-\mathcal{W}[\mathcal{J}]+\int_{t,\mathbf{x}}\mathcal{J}_{i}\Psi_{i}\right]\,,\qquad\hbox{where}\quad\Psi_{i}\equiv\langle\Phi_{i}\rangle_{\cal J}=\dfrac{\delta{\cal W}}{\delta{\cal J}_{i}}\,$, (12)
with $\Phi=(\mathbf{v},\bar{\mathbf{v}},\pi,\bar{\pi})$ denoting the multiplet of fields, $\mathcal{J}=(\mathbf{j},\bar{\mathbf{j}},K,\bar{K})$ the multiplet of associated sources, and $\Psi=(\mathbf{u},\bar{\mathbf{u}},\Pi,\bar{\Pi})$ the multiplet of expectation values of the fields in the presence of the sources $\cal J$}. One can obtain from $\Gamma$ the vertex functions, defined as
$\Gamma^{(m,n)}_{\alpha_{1}...\alpha_{m+n}}(t_{1},\mathbf{x}_{1};\ldots;t_{m+n},\mathbf{x}_{m+n})\equiv\dfrac{\delta^{m+n}\Gamma}{\delta u_{\alpha_{1}}(t_{1},\mathbf{x}_{1})\ldots\delta\bar{u}_{\alpha_{m+n}}(t_{m+n},\mathbf{x}_{m+n})}\Big|_{\mathbf{u},\bar{\mathbf{u}}=0}\,$, (13)
where again the $m$ first derivatives are with respect to $\mathbf{u}$ fields and the $n$ last with respect to $\bar{\mathbf{u}}$ fields. The knowledge of the set of $G^{(m,n)}$, or equivalently of the set of $\Gamma^{(m,n)}$, provides the statistical theory for the model under study.
Let us now briefly explain how the symmetries lead to exact relations between these functions, which are called Ward identities, and which can be very efficiently exploited within the FRG formalism. In fact, it turns out to be particularly useful to consider extended symmetries, which correspond to continuous infinitesimal transformations of coordinates and fields which do not leave the action strictly invariant, but yields a variation at most linear in the fields. One can indeed derive from such extended symmetries exact Ward identities, which deeply constrain the set of $G^{(m,n)}$ (or equivalently the set of $\Gamma^{(m,n)}$[45].
This procedure has been expounded in details, in eg. Refs. [38, 39] for the Navier-Stokes equation, or in eg. Refs. [46, 28] for the Burgers–KPZ equation. Hence, we only provide here a summary, and refer the interested reader to these references for technical details. The first noticeable extended symmetries are shifts of the pressure fields, $\pi(t,\mathbf{x})\to\pi(t,\mathbf{x})+\epsilon(t,\mathbf{x})$, and $\bar{\pi}(t,\mathbf{x})\to\bar{\pi}(t,\mathbf{x})+\bar{\epsilon}(t,\mathbf{x})$, with $\epsilon(t,\mathbf{x})$ and $\bar{\epsilon}(t,\mathbf{x})$ infinitesimal scalar functions of time and coordinates. The corresponding Ward identities impose that the variation of the effective action $\Gamma$ under these transformations is equal to that of the action $\cal S$}, which entails that the whole pressure sector is not renormalised. A second extended symmetry is the time-dependent Galilean symmetry, which corresponds to the transformation
@@
with $\epsilon(t)$ an infinitesimal vectorial function of time. The corresponding Ward identities exactly relate a vertex function $\Gamma^{(m+1,n)}$ with one vanishing momentum among the first $m+1$ (ie. associated with a velocity) to lower order ones $\Gamma^{(m,n)}$. We here simply give the two lowest-order identities
@@
with $\bar{\epsilon}(t)$ an infinitesimal vectorial function of time. The corresponding Ward identities exactly impose that a vertex function $\Gamma^{(m,n)}$ with one vanishing momentum among the last $n$ (i.e associated with a response velocity) vanishes, except the two lowest-order ones, which are fixed to
@@
This symmetry also yields constraints on the vertex functions, in particular